# From Small to Large: A Graph Convolutional Network Approach for Solving Assortment Optimization Problems

### v1

@@ Abstract
Assortment optimization involves selecting a subset of substitutable products (subject to certain constraints) to maximize the expected revenue. It is a classic problem in revenue management and finds applications across various industries. However, the problem is usually NP-hard due to its combinatorial and non-linear nature. In this work, we explore how graph concolutional networks (GCNs) can be leveraged to efficiently solve constrained assortment optimization under the mixed multinomial logit choice model. We first develop a graph representation of the assortment problem, then train a GCN to learn the patterns of optimal assortments, and lastly propose two inference policies based on the GCN’s output. Due to the GCN’s inherent ability to generalize across inputs of varying sizes, we can use a GCN trained on small-scale instances to facilitate large-scale instances. Extensive numerical experiments demonstrate that given a GCN trained on small-scale instances (e.g., with 20 products), the proposed policies can achieve superior performance (90%+ optimality) on large-scale instances (with up to 2,000 products) within seconds, which outperform existing heuristic policies in both performance and efficiency. Furthermore, we extend our framework to a model-free setting where the underlying choice model is unknown but transaction data is available. We also conduct numerical experiments to demonstrate the effectiveness and efficiency of our proposed policies in this setting.
@@ 1 Introduction
Given the probabilities, we still need inference policies to derive solutions. In this work, we first propose an index policy that sorts all products in a descending order by their indices and then evaluates all feasible assortments consisting of the top-$n$ products according to the indices for $n=1,2,\dots,N$, where $N$ is the number of products. Such an approach is inspired by the well-studied “revenue-order policy” in the literature, which can be interpreted as an index policy based on prices. In our work, we implement the index policy based on the probabilities predicted by the GCN, which is referred to as the “GCN” policy. This policy is computationally efficient because the number of evaluated assortments is no greater than $N$, which is significantly smaller than the total number of assortments $12^{N}-$ 1. Moreover, we also consider a “GCN-LS” policy that adopts the local-search technique in Gallego et al. (2024) to further improve the solution derived by the GCN policy.
@@ 2 Model
In this work, we study the assortment optimization problem under linear constraints. Consider $N$ products with the fixed price vector $\boldsymbol{r}\in\mathbb{R}_{+}^{N}$. Without loss of generality, we assume that the product costs are zero. Let $[N]$ denote the whole set $\{1,2,\dots,N\}$ and 00 denote the no-purchase option. Given an assortment $S\subseteq[N]$ represented by $\boldsymbol{x}_{S}=\left\{x_{S,j}=\mathbbm{1}_{\{j\in S\}}\right\}_{j=1}^{N}$, the choice probability of product $j$ is denoted by $\phi_{j}(\boldsymbol{x}_{S})$, which can be captured by some choice model. Moreover, we consider $M$ constraints represented by $\boldsymbol{A}\boldsymbol{x}\leq\boldsymbol{b}$, where $\boldsymbol{A}\in\mathbb{R}^{M\times N}$ and $\boldsymbol{b}\in\mathbb{R}^{M\times 1}$. Then, the constrained assortment optimization problem can be formulated as a non-linear integer program as follows:
@@ 3.2 GCN
    ($\mathcal{F}\left(\{\boldsymbol{m}_{1},\boldsymbol{m}_{2},\dots,\boldsymbol{m}_{S}\}\right):=\boldsymbol{m}^{\prime}\in\mathbb{R}^{1\times d_{\text{mid}}},\text{ where $m^{\prime}_{\ell}=\sum_{s=1}^{S}\frac{e^{m_{s,\ell}}}{\sum_{s^{\prime}=1}^{S}e^{m_{s^{\prime},\ell}}}\times m_{s,\ell}$}.$
@@
    $\boldsymbol{h}^{S}_{k}$ :=($:=\mathcal{F}\left(\left\{\boldsymbol{m}^{2}_{jk}\right\}_{j=1}^{N}\right)+\boldsymbol{l}_{k}^{2}\in\mathbb{R}^{1\times d_{\text{mid}}}$
@@
    $\boldsymbol{h}^{C}_{i}$ :=($:=\mathcal{F}\left(\left\{\boldsymbol{m}^{3}_{ji}\right\}_{j=1}^{N}\right)+\boldsymbol{l}^{3}_{i}\in\mathbb{R}^{1\times d_{\text{mid}}}.$
@@ 4 Numerical Experiments
    CONIC: We implement the conic programming approach in Şen et al. (2018), which is significantly faster than the traditional MILP formulation in Méndez-Díaz et al. (2014). We set the MIPGap threshold as $0.1\%$ and the time limit of the conic program as 600 seconds.
@@
101010102020202030303030404040405050505060606060%$80$%$100$%$N_{\text{train}}$ompetitive Ratio
10010010010020020020020030030030030040040040040050050050050060606060%$80$%$100$%$N_{\text{test}}$ompetitive Ratio
555510101010151515152020202060606060%$80$%$100$%$K_{\text{test}}$ompetitive Ratio
222244446666888860606060%$80$%$100$%$M_{\text{test}}$ompetitive Ratio
@@ Appendix B Discussion of Other Choice Models
Then, we provide the graph representation under the exponomial model. In Figure 5, we use the same graph representation in Figure 1 with $K=1$. However, the feature vectors are slightly different. The feature vector of the only customer node is $[0,1,0]$; the feature vector of product node $j$ is $\left[r_{j},0,1\right]$; the feature vector of constraint node $i$ is $[0,0,b_{i}]$. The feature vector of the edge from the customer node to product node $j$ is $[v_{j},\lambda_{j},0]$, and the feature vector of the edge between product node $j$ and constraint node $i$ is $[0,0,A_{ij}]$.
@@
As for the graph representation, we need to modify Figure 1 to include nest nodes ($\bigstar$). In Figure 6, we provide the graph representation of the problem under the NL model. Note that when connecting nest nodes and product nodes, we only connect each nest node with products in the nest. The feature vector of the only customer node is $[0,0,1,0]$; the feature vector of product node $j$ is $\left[r_{j},0,0,1\right]$; the feature vector of constraint node $i$ is $[0,0,0,b_{i}]$; the feature vector of nest node $h$ is $[0,\gamma_{h},0,0]$. The feature vector of the edge from the customer node to nest node $h$ is $[1,0,0]$, the feature vector of the edge between nest node $h$ and product node $j$ is $[0,v_{hj},0]$, and the feature vector of the edge between product node $j$ and constraint node $i$ is $[0,0,A_{ij}]$.
@@
As for the graph representation, in addition to edges in Figure 1, we introduce the product-to-product edges to represent the transition probabilities. In Figure 7, we provide the graph representation of the problem under the MC model. The feature vector of the only customer node is $[0,1,0]$; the feature vector of product node $j$ is $\left[r_{j},0,1\right]$; the feature vector of constraint node $i$ is $[0,0,b_{i}]$. The feature vector of the edge from the customer node to product node $j$ is $[\beta_{j},0,0]$, the feature vector of the directed edge from product node $j_{1}$ to product node $j_{2}$ is $[0,\rho_{j_{1}j_{2}},0]$, and the feature vector of the edge between product node $j$ and constraint node $i$ is $[0,0,A_{ij}]$.

### v2

@@ Abstract
Assortment optimization involves selecting a subset of substitutable products (subject to certain constraints) to maximize the expected revenue. It is a classic problem in revenue management and finds applications across various industries. However, the problem is usually NP-hard due to its combinatorial and non-linear nature. In this work, we explore how graph convolutional networks (GCNs) can be leveraged to efficiently solve constrained assortment optimization under the mixed multinomial logit choice model. We first develop a graph representation of the assortment problem, then train a GCN to learn the patterns of optimal assortments, and lastly propose two inference policies based on the GCN’s output. Due to the GCN’s inherent ability to generalize across inputs of varying sizes, we can use a GCN trained on small-scale instances to facilitate large-scale instances. Extensive numerical experiments demonstrate that given a GCN trained on small-scale instances (e.g., with 20 products), the proposed policies can achieve superior performance (90%+ optimality) on large-scale instances (with up to 2,000 products) within seconds, which outperform existing heuristic policies in both performance and efficiency. Furthermore, we extend our framework to a model-free setting where the underlying choice model is unknown but transaction data is available. We also conduct numerical experiments to demonstrate the effectiveness and efficiency of our proposed policies in this setting.
@@ 1 Introduction
Given the probabilities, we still need inference policies to derive solutions. In this work, we first propose an index policy that sorts all products in a descending order by their indices and then evaluates all feasible assortments consisting of the top-$n$ products according to the indices for $n=1,2,\dots,N$, where $N$ is the number of products. Such an approach is inspired by the well-studied “revenue-order policy” in the literature, which can be interpreted as an index policy based on prices. In our work, we implement the index policy based on the probabilities predicted by the GCN, which is referred to as the “GCN” policy. This policy is computationally efficient because the number of evaluated assortments is no greater than $N$, which is significantly smaller than the total number of assortments $2^{N}-1$. Moreover, we also consider a “GCN-LS” policy that adopts the local-search technique in Gallego et al. (2024) to further improve the solution derived by the GCN policy.
@@ 2 Model
In this work, we study the assortment optimization problem under linear constraints. Consider $N$ products with the fixed price vector $\boldsymbol{r}\in\mathbb{R}_{+}^{N}$. Without loss of generality, we assume that the product costs are zero. Let $[N]$ denote the whole set $\{1,2,\dots,N\}$ and 0 denote the no-purchase option. Given an assortment $S\subseteq[N]$ represented by $\boldsymbol{x}_{S}=\left\{x_{S,j}=\mathbbm{1}_{\{j\in S\}}\right\}_{j=1}^{N}$, the choice probability of product $j$ is denoted by $\phi_{j}(\boldsymbol{x}_{S})$, which can be captured by some choice model. Moreover, we consider $M$ constraints represented by $\boldsymbol{A}\boldsymbol{x}\leq\boldsymbol{b}$, where $\boldsymbol{A}\in\mathbb{R}^{M\times N}$ and $\boldsymbol{b}\in\mathbb{R}^{M\times 1}$. Then, the constrained assortment optimization problem can be formulated as a non-linear integer program as follows:
@@ 3.2 GCN
    $\mathcal{F}\left(\{\boldsymbol{m}_{1},\boldsymbol{m}_{2},\dots,\boldsymbol{m}_{S}\}\right):=\boldsymbol{m}^{\prime}\in\mathbb{R}^{1\times d_{\text{mid}}},\text{ where $m^{\prime}_{\ell}=\sum_{s=1}^{S}\frac{e^{m_{s,\ell}}}{\sum_{s^{\prime}=1}^{S}e^{m_{s^{\prime},\ell}}}\times m_{s,\ell}$}.$
@@
    $\boldsymbol{h}^{S}_{k}$ $:=\mathcal{F}\left(\left\{\boldsymbol{m}^{2}_{jk}\right\}_{j=1}^{N}\right)+\boldsymbol{l}_{k}^{2}\in\mathbb{R}^{1\times d_{\text{mid}}}$
@@
    $\boldsymbol{h}^{C}_{i}$ $:=\mathcal{F}\left(\left\{\boldsymbol{m}^{3}_{ji}\right\}_{j=1}^{N}\right)+\boldsymbol{l}^{3}_{i}\in\mathbb{R}^{1\times d_{\text{mid}}}.$
@@ 4 Numerical Experiments
    CONIC: We implement the conic programming approach in Şen et al. (2018), which is significantly faster than the traditional MILP formulation in Méndez-D´ıaz et al. (2014). We set the MIPGap threshold as $0.1\%$ and the time limit of the conic program as 600 seconds.
@@
[…]
@@ Appendix B Discussion of Other Choice Models
Then, we provide the graph representation under the exponomial model. In Figure 5, we use the same graph representation in Figure 1 with $K=1$. However, the feature vectors are slightly different. The feature vector of the only customer node is $[0,1,0]$; the feature vector of product node $j$ is $\left[r_{j},0,1\right]$; the feature vector of constraint node $i$ is $0,0,b_{i}]$. The feature vector of the edge from the customer node to product node $j$ is $v_{j},\lambda_{j},0]$, and the feature vector of the edge between product node $j$ and constraint node $i$ is $0,0,A_{ij}]$.
@@
As for the graph representation, we need to modify Figure 1 to include nest nodes ($\bigstar$). In Figure 6, we provide the graph representation of the problem under the NL model. Note that when connecting nest nodes and product nodes, we only connect each nest node with products in the nest. The feature vector of the only customer node is $[0,0,1,0]$; the feature vector of product node $j$ is $\left[r_{j},0,0,1\right]$; the feature vector of constraint node $i$ is $0,0,0,b_{i}]$; the feature vector of nest node $h$ is $0,\gamma_{h},0,0]$. The feature vector of the edge from the customer node to nest node $h$ is $[1,0,0]$, the feature vector of the edge between nest node $h$ and product node $j$ is $0,v_{hj},0]$, and the feature vector of the edge between product node $j$ and constraint node $i$ is $0,0,A_{ij}]$.
@@
As for the graph representation, in addition to edges in Figure 1, we introduce the product-to-product edges to represent the transition probabilities. In Figure 7, we provide the graph representation of the problem under the MC model. The feature vector of the only customer node is $[0,1,0]$; the feature vector of product node $j$ is $\left[r_{j},0,1\right]$; the feature vector of constraint node $i$ is $0,0,b_{i}]$. The feature vector of the edge from the customer node to product node $j$ is $\beta_{j},0,0]$, the feature vector of the directed edge from product node $j_{1}$ to product node $j_{2}$ is $0,\rho_{j_{1}j_{2}},0]$, and the feature vector of the edge between product node $j$ and constraint node $i$ is $0,0,A_{ij}]$.